OpenStudy (anonymous):

I keep getting this wrong, can I please get some help xD Trying to estimate the area under the graph f(x) = 7 + x^2 using six approximating rectangles and left endpoints (L6).

OpenStudy (anonymous):

I've already calculated R3, R6, L3, M3, & M6 but I keep getting L6 wrong.

ganeshie8 (ganeshie8):

whats the interval ?

OpenStudy (anonymous):

Should be 0.5 right?

OpenStudy (anonymous):

Cause it says to use 6 rectangles

OpenStudy (anonymous):

Ohhh I didn't give you the whole question! Forgive me haha! Here it is... Estimate the area under the graph of f(x)=7+x2 from x=−1 to x=2

OpenStudy (anonymous):

I get this... \[L _{6}= (7/2)+(7.25/2)+(7/2) +(7.25/2)+(7/2)+(1/2)(7+1.5^2)\] But apparently I'm doing something wrong :/

OpenStudy (anonymous):

O.o haha i don't get it

ganeshie8 (ganeshie8):

width of each interval = (2--1)/6 = 3/6 = 1/2 this is correct in your work

OpenStudy (anonymous):

Ok :) next? What part of my calculations is wrong?

ganeshie8 (ganeshie8):

next, you're splitting the given interval (-1, 2) into 6 subinterval sand evaluating the f(x) at the left end point in each subinterval

OpenStudy (anonymous):

Right

ganeshie8 (ganeshie8):

f(x) = 7 + x^2 the first subinterval starts at -1, so the height of first rectangle is 7 + (-1)^2 = 7+1 = 8 yes ?

OpenStudy (anonymous):

Yep thats right!

OpenStudy (anonymous):

Ohhhh >.<

OpenStudy (anonymous):

I said 7, didn't I? -.-

ganeshie8 (ganeshie8):

the second subinterval starts at x = -1 + 1/2 = -1/2 so you need to evaluate f(-1/2) for the height of second subinterval

OpenStudy (anonymous):

Right!!

ganeshie8 (ganeshie8):

so area = 1/2*[f(-1) + f(-1/2) + f(0) + f(1/2) + f(1) + f(3/2)]

OpenStudy (anonymous):

Exactly! Thats true! Such a silly mistake :D lol

OpenStudy (anonymous):

But I'm still getting it wrong :/

ganeshie8 (ganeshie8):

try entering 23.375

ganeshie8 (ganeshie8):

or 187/8

OpenStudy (anonymous):

Thats right! :D

OpenStudy (anonymous):

Thanks so much!!! <3

ganeshie8 (ganeshie8):

yw!

OpenStudy (anonymous):

:)